OpenStudy (anonymous):

If tan x = .88, what is cot x?

OpenStudy (anonymous):

will someone explain

OpenStudy (solomonzelman):

tan and cot are multiplicative inverses of each other.

OpenStudy (anonymous):

@lskjdflsdjf

OpenStudy (solomonzelman):

Hello? is tan(x)=.88 then cot(x)=1/.88=100/88=25/22

OpenStudy (anonymous):

what u mena

OpenStudy (solomonzelman):

I am just applying the law that states. \[\tan(x)=\frac{1}{\cot(x)}\]

OpenStudy (anonymous):

ok. so what do i do ?

OpenStudy (solomonzelman):

plug in .88 for x, and say that if tan(x)=.88, then cot(x)=1 / .88 now simplify 1 / .88

OpenStudy (anonymous):

How do i simplify that

OpenStudy (solomonzelman):

\[\frac{1}{.88}=\frac{100}{88}==\frac{25}{22}=1\frac{3}{22}=1.~136~136~136....≈1.14\]